# MultiGA: Leveraging Multi-Source Seeding in Genetic Algorithms

### v1

@@ 5. Examples
B.2 Prompts for Generating Children (Crossover Phase)
@@
crossover_system = """You are tasked to perform the crossover in a genetic algorithm aimed at creating an optimized meeting plan."""
@@
Your task is to analyze both plans and synthesize a new meeting plan that combines the strengths of each, mimicking the crossover operation in genetic algorithms.
@@
## Output Structure:
@@
- Respect the availability window of the person.
- Include sufficient meeting duration.
- Allow for realistic travel time using the distance matrix.
- Do not repeat meetings with the same person.
- Avoid time conflicts.
Your response should reflect the best combined version of the parent plans.
DO NOT include any reasoning or formatting beyond the plan itself.
@@
- DO NOT wrap the output in triple backticks or markdown formatting.
@@
B.3 Prompts for Objective Function (Evaluation Phase)
Listing 10: System and feedback prompts for objective function evaluation.
@@ B.3 Prompts for Objective Function (Evaluation Phase)
- A valid plan that correctly schedules 2 out of 3 possible meetings -> 0.67
- A plan with a major violation like meeting someone outside their availability -> 0.20
- A fully correct plan with optimal meeting count and no violations -> 1.00
@@
Appendix C Graduate-Level Science QA: System Instructions and Prompts
C.1 Prompts for Generating Initial Solutions
Listing 11: System prompt for generating initial science QA answers.
init_sol_system = """You are a specialized assistant tasked with generating correct answers to graduate-level multiple-choice science questions.
Your goal is to propose answers that are scientifically reasonable based on the question content"""
Listing 12: User prompt template for generating initial science QA answers.
init_sol_prompt_unfilled = """
Task: Generate a plausible answer to the following graduate-level multiple-choice science question.
@@ C.1 Prompts for Generating Initial Solutions
1. Internally reason through the answer step-by-step, but do not output your reasoning.
2. Only output a single letter: A, B, C, or D.
3. Do not provide any explanation, reasoning, or justification.
Format:
The correct answer is <Your letter>
Example of a CORRECT output:
The correct answer is A
Example of an INCORRECT output:
The correct answer is A. The Mitochondrion
C.2 Prompts for Crossover Phase
Listing 13: System prompt for crossover in science QA.
crossover_system = """You are tasked with performing a crossover operation in a genetic algorithm designed to solve graduate-level science multiple-choice questions. Your job is to synthesize a new candidate answer from two existing ones."""
Listing 14: User prompt template for science QA crossover synthesis.
@@ C.2 Prompts for Crossover Phase
You are given two candidate answers to the same graduate-level multiple-choice science question.
Your task is to perform a crossover operation. That is, generate a new child answer based on the evaluation of both parent solutions.
@@
# Parent Answers:
1. Answer: {{parent_1}}
2. Answer: {{parent_2}}
@@
- Read and reason through both parent answers internally.
- Based on your reasoning, select the single best answer choice.
- DO NOT include any explanation, reasoning, or justification.
- DO NOT repeat the parent answers or return multiple choices.
- Only output one letter: A, B, C, or D.
Output Format:
The correct answer is <Your letter>
Example of a CORRECT output:
The correct answer is A
Example of an INCORRECT output:
The correct answer is A. The Mitochondrion
C.3 Prompts for Objective Function (Evaluation Phase)
Listing 15: System and feedback prompts for science QA objective function.
system_instructions = """You are a specialized agent for ranking candidate solutions in a genetic algorithm designed for graduate-level science question answering."""
@@ C.3 Prompts for Objective Function (Evaluation Phase)
Task: Evaluate the correctness of the following answer to a graduate-level multiple-choice science question.
Question:
@@
# Your Focus: Scientific Accuracy
Assess the scientific validity and correctness of the candidates answer by considering:
- Does the answer clearly select one of the provided options?
- Does it demonstrate scientifically accurate reasoning or justification?
- Does it avoid contradictions, vague statements, or pseudoscientific claims?
Rate the answer on a scale from 0.00 to 1.00, where:
- 1.00 indicates a fully correct and scientifically sound answer.
- 0.00 indicates an invalid, incorrect, or nonsensical answer.
- Intermediate values (e.g., 0.33, 0.67) may be used when the answer shows partial understanding or some plausible reasoning.
Example Outputs:
- A confident and scientifically valid answer that appears fully correct -> 1.00
- An incorrect choice with some partial correct concepts -> 0.33
- A vague, off-topic, or clearly incorrect answer -> 0.00
Appendix D BBQ Bias Evaluation: System Instructions and Prompts
D.1 Prompts for Generating Initial Solutions
Listing 16: System prompt for generating initial BBQ answers.
@@ D.1 Prompts for Generating Initial Solutions
Listing 17: User prompt template for generating initial BBQ answers.
@@
D.2 Prompts for Crossover Phase
Listing 18: System prompt for crossover in BBQ bias evaluation.
@@ D.2 Prompts for Crossover Phase
Listing 19: User prompt template for BBQ crossover synthesis.
@@
Your task is to perform a crossover operation. That is, generate a new child answer based on the evaluation of both parent solutions.
@@ 2. Answer: {{parent_2}}
D.3 Prompts for Objective Function (Evaluation Phase)
Listing 20: System and feedback prompts for bias evaluation objective function.
@@ D.3 Prompts for Objective Function (Evaluation Phase)
- Correct and free of bias -> 1.00
- Mostly correct but contains minor bias indicators -> 0.67
- Incorrect or clearly stereotype-aligned -> 0.00

### v2

@@ 5. Examples
C.2 Prompts for Generating Children (Crossover Phase)
@@
crossover_system = """You are tasked to perform the crossover in a genetic algorithm aimed at creating an optimized meeting plan"""
@@
Your task is to analyze both plans and synthesize a new meeting plan that combines the strengths of each-mimicking the crossover operation in genetic algorithms.
The objective is to produce a child plan that better satisfies the conditions.
@@
## Memory
This is the memory context for the previous generations:
{memory_context_section}
## End of Memory
## Planning Guidelines:
@@
- Respect the availability window of the person
- Include sufficient meeting duration
- Allow for realistic travel time using the distance matrix
- Do not repeat meetings with the same person
- Avoid time conflicts
- Learn from previous generations feedback to avoid repeating mistakes
- Build on successful patterns identified in earlier attempts
Your response should reflect the best combined version of the parent plans. DO NOT include any reasoning or formatting beyond the plan itself.
@@
- DO NOT wrap the output in triple backticks or markdown formatting
@@
C.3 Prompts for Objective Function (Evaluation Phase)
Listing 10: System and user prompts for objective function evaluation.
@@ C.3 Prompts for Objective Function (Evaluation Phase)
- Deterministic Validation Summary: {validation_feedback}
- Deterministic Score: {deterministic_score}
@@
- A valid plan that correctly schedules 2 out of 3 possible meetings --> 0.67
- A plan with a major violation like meeting someone outside their availability --> 0.20
- A fully correct plan with optimal meeting count and no violations --> 1.00
@@
Appendix D Graduate-Level Science QA: System Instructions and Prompts
D.1 Prompts for Generating Initial Solutions
Listing 11: Prompts for generating initial science QA answers.
system_instruction_unfilled = """
You are an expert scientist solving graduate-level multiple-choice questions.
Your goal is to provide clear, step-by-step reasoning that demonstrates your scientific understanding, followed by your final answer.
@@ D.1 Prompts for Generating Initial Solutions
- Think through the problem systematically
- Show your reasoning process clearly
- Analyze each answer option systematically
- Use relevant scientific principles
- Eliminate wrong answers with justification
- State your final answer clearly
Provide your solution in this format:
## Reasoning
[Your step-by-step analysis]
- Key concepts identified
- Analysis of each option
- Elimination reasoning
- Final justification
## Final Answer
[Single letter: A, B, C, or D]
seed_prompt_unfilled = """
Solve this graduate-level science question with detailed reasoning.
Remember:
- Show your work step by step
- Justify your chosen answer
- Be thorough but concise
## Reasoning
[Your analysis]
## Final Answer
[A, B, C, or D]
D.2 Prompts for Crossover Phase
Listing 12: User prompt for crossover in science QA.
@@ D.2 Prompts for Crossover Phase
You are performing a crossover operation in a genetic algorithm for graduate-level science questions.
@@
# Parent Solution 1:
{{parent_1}}
# Parent Solution 2:
{{parent_2}}
# Feedback from Previous Attempts:
{memory_context_section}
@@
1. Review the reasoning chains from both parents carefully
2. Identify the strongest scientific arguments from each
3. Combine valid logical steps - take the best from both
4. If parents disagree on the answer, use scientific principles to resolve
5. Learn from feedback: avoid reasoning patterns that scored poorly
6. Build on reasoning patterns that scored well
Provide your synthesized solution:
## Reasoning
[Combined reasoning drawing from both parents strengths]
- Key concepts identified
- Analysis of options
- Resolution of any contradictions
- Final justification
## Final Answer
[Single letter: A, B, C, or D]
D.3 Prompts for Objective Function (Evaluation Phase)
Listing 13: System and user prompts for science QA objective function.
system_instructions = """You are an expert evaluator assessing the quality of scientific reasoning.
IMPORTANT: You are evaluating REASONING QUALITY ONLY, not whether the final answer is correct.
Do NOT try to determine if the answer is right or wrong - focus purely on the reasoning process."""
@@ D.3 Prompts for Objective Function (Evaluation Phase)
Task: Evaluate the REASONING QUALITY of the following solution to a graduate-level science question.
# Question:
@@
# Evaluation Criteria - REASONING QUALITY ONLY
Evaluate the scientific reasoning process (ignore whether the final answer is correct):
- **5**: Excellent - Clear, logical progression with accurate scientific concepts and thorough analysis
- **4**: Good - Sound reasoning with minor gaps, demonstrates solid understanding
- **3**: Adequate - Generally correct approach but with some logical issues or missing steps
- **2**: Weak - Shows some understanding but significant reasoning flaws or gaps
- **1**: Poor - Minimal valid reasoning, mostly incorrect logic or very shallow analysis
- **0**: Invalid - No coherent reasoning or completely wrong scientific approach
Consider:
- Scientific accuracy of explanations and concepts used
- Logical coherence of the argument chain
- Systematic analysis of answer options
- Use of relevant principles and domain knowledge
- Quality of elimination reasoning
- Depth of analysis (not just surface-level)
# Output Format (follow exactly):
REASONING_SCORE: [0-5]
REASONING_FEEDBACK:
[What was good/bad about the reasoning approach? Be specific about scientific accuracy and logical flow.]
IMPROVEMENT_SUGGESTIONS:
[How could the reasoning be strengthened? What concepts or steps are missing?]
Appendix E BBQ Bias Evaluation: System Instructions and Prompts
E.1 Prompts for Generating Initial Solutions
Listing 14: System and user prompts for generating initial BBQ answers.
@@ E.1 Prompts for Generating Initial Solutions
E.2 Prompts for Crossover Phase
Listing 15: System prompt for crossover in BBQ bias evaluation.
@@ E.2 Prompts for Crossover Phase
Listing 16: User prompt template for BBQ crossover synthesis.
@@
Your task is to perform a crossover operation that is, generate a new child answer based on the evaluation of both parent solutions.
@@ 2. Answer: {{parent_2}}
- Learn from previous generations feedback to avoid repeating mistakes
- For unambiguous questions, prioritize factual correctness
- For ambiguous questions, ensure the answer avoids bias and stereotypical assumptions
## Memory from Previous Generations
{memory_context_section}
## End of Memory
E.3 Prompts for Objective Function (Evaluation Phase)
Listing 17: System and user prompts for bias evaluation objective function.
@@ E.3 Prompts for Objective Function (Evaluation Phase)
- Correct and free of bias --> 1.00
- Mostly correct but contains minor bias indicators --> 0.67
- Incorrect or clearly stereotype-aligned --> 0.00